# Probing Neural Topology of Large Language Models

### v2

@@ 4 Topological analysis and causal intervention
Figure 4: (a-b) Occurrence frequency of hub nodes in (a) Qwen2.5-0.5B and (b) Qwen2.5-1.5B on MMLU benchmark. (c) Accuracy on MMLU benchmark of Qwen2.5-0.5B and Qwen2.5-1.5B under different interventions of top 1% neurons.
@@
To determine if LLMs rely on their neural topology, we conducted a series of interventional experiments. We disabled a selected 1% of neurons in the middle layer of Qwen2.5-0.5B and Qwen2.5-1.5B by pinning their activations to zero at all tokens and measured the impact on the MMLU benchmark (Hendrycks et al., 2021). We compared three distinct neuron selection strategies: random selection, selecting neurons with the highest activation, and selecting neurons with the highest topological degree. As shown in the Figure 4(c), the results demonstrate a clear hierarchy of neuronal importance. While disabling random neurons caused a negligible accuracy drop (<1.0%), targeting neurons by either activation or topology incurred a substantial performance collapse of at least 20%. Critically, the topology-based intervention was the most detrimental. Disabling the top 1% of hub neurons resulted in a catastrophic accuracy drop of 45.4% and 57.5% for the two tested models, which is over 20% worse than even the activation-based strategy. These experiments provide causal evidence that LLMs actively utilize their underlying topological structure, particularly their hub neurons, for computation.
Our analysis also revealed a notable phenomenon of functional specialization (Cole et al., 2014), where neural topology displays distinct patterns for different subjects. We found a striking example by identifying certain neuron that consistently specializes in STEM subjects, while remaining almost entirely dormant when processing texts from the social science subjects. This suggests a subject-based organization within LLMs, and the details are provided in Appendix A.5.
@@ 5 Applications
In this section, we explore the practical implications of neural topology through three proof-of-concept applications: model pruning for efficiency, hallucination detection for reliability, and model fingerprinting for security. Though preliminary in their current forms, these case studies showcase the versatile utility of neural topology and lay the groundwork for future work to translate these concepts into robust, system-level solutions.
5.1 Model pruning
@@
The finding that high-degree neurons are functionally dominant (Section 4) naturally suggests a pruning strategy: disabling low-degree neurons to reduce computational cost. To test this, we prune these neurons by zeroing out their activations during inference. We apply this technique to a middle layer of Qwen2.5-0.5B and Qwen2.5-1.5B and evaluate the performance on the MMLU benchmark (Hendrycks et al., 2021). The results, presented in Figure 5(a), confirm the models’ robustness. Performance degradation is minimal even with substantial pruning; for the 1.5B model, accuracy drops by just 2.86% when 10% of neurons are pruned, and by 11.73% when 25% are removed. Strikingly, the models maintain above-random-guess accuracy until half of the low-degree neurons are pruned. These results highlight the potential for developing more sophisticated pruning methods based on neural topology.
5.2 Hallucination detection
@@
Building on our finding that subtle topological patterns in LLMs can predict text generation performance, we now investigate if these patterns can also detect a critical failure mode: hallucination, analogous to identifying deception from human brain signals. To create distinct genuine and hallucinating states, we construct inputs from the TruthfulQA dataset (Lin et al., 2022) by concatenating each of its 817 question with its corresponding true and false answers, respectively, resulting in a dataset of 5918 samples (true/false classification). We then extract the neural topology from the LLM as it processes these inputs (Equations 1-3). For this task, we adapt probing from regression to binary classification by modifying the linear/MLP probe layer to have two output channels ($\mathbf{W}_{1}\in\mathbb{R}^{2\times(n\times n)}$ and $\mathbf{W}_{3}\in\mathbb{R}^{2\times d}$ in Equations (4-5)) and replacing the MSE loss with a cross-entropy loss:
$\textbf{Loss}:$~ $\mathcal{L}(\mathbf{X})=$ $\frac{1}{N}\sum_{i=1}^{N}\mathtt{CROSS\_ENTROPY}(\hat{y}_{i},y_{i}),$ (8)
@@
where $\hat{y}$ and $y$ are the prediction and ground-truth for hallucination (0/1). We compare against linear and MLP probes trained on the neural activation at the last-token position.
@@
We split the dataset into training and test sets using an 8:2 ratio to evaluate our hallucination detection framework. As shown in Figure 5(b), probing neural topology substantially outperforms probing neural activation for all three tested models (GPT-2, Pythia-160M, and Qwen2.5-0.5B), with accuracy gains of up to 9.73%. This superior performance suggests that distinct topological patterns emerge when an LLM is generating a factual response versus hallucinating. To validate this hypothesis directly, we introduce a neural topology coupling index,
$C_{TT}=\mathtt{AVG}(\{\rho(A_{i},A_{j})|A_{i},A_{j}\in A_{T}\}),C_{HH}=\mathtt{AVG}(\{\rho(A_{i},A_{j})|A_{i},A_{j}\in A_{H}\}),$ (9)
$C_{TH}=\mathtt{AVG}(\{\rho(A_{i},A_{j})|A_{i}\in A_{T},A_{j}\in A_{H}\}),$ (10)
@@
Figure 5: (a) Accuracy on MMLU benchmark under different levels of model pruning. (b) Accuracy of hallucination detection probed from neuron activation and topology on TruthfulQA dataset. (c) Coupling index of neural topology for hallucination on TruthfulQA datset.
5.3 Model matching and fingerprinting
In this section, we further investigate potential topological similarity across different LLMs to answer the following question: can we detect the genetic relationships of LLMs (e.g. same model family, or finetuning) from their internal neural topology? To this end, we extend graph probing with contrastive learning to perform graph matching, as illustrated in Figure 6. Specifically, suppose we feed a batch of $B$ token sequences into two LLMs, $\Omega$ and $\Gamma$. We compute the corresponding neural connectivity graphs and use two GNN probes to encode them into representations $\mathbf{Z}_{\Omega}=[\mathbf{z}^{\Omega}_{1},\ldots,\mathbf{z}^{\Omega}_{B}]$ and $\mathbf{Z}_{\Gamma}=[\mathbf{z}^{\Gamma}_{1},\ldots,\mathbf{z}^{\Gamma}_{B}]$. Matching is implemented using a contrastive cross-entropy loss that encourages alignment between graph representations from the same input texts:
$\mathcal{L}=\sum_{i=1}^{B}{\mathtt{CROSS\_ENTROPY}(\mathcal{S}_{i,:},\mathcal{T}_{i,:})}+\sum_{j=1}^{B}{\mathtt{CROSS\_ENTROPY}(\mathcal{S}_{:,j},\mathcal{T}_{:,j})},$ (13)
where $\mathcal{S}$ is the similarity matrix by taking inner product of graph representations and $\mathtt{IDENTITY}$ indicates identity matrix which is the target $\mathcal{T}$ for graph matching. After training the graph probes contrastively on a shared set of training texts, the out-of-sample graph matching performance serves as an indicator of neural topology similarity between two LLMs. To evaluate this, we adopt the commonly used GAUC metric (Li et al., 2019) (see Appendix A.6 for details).
Matching LLM $\mathbf{\Omega}$ LLM $\mathbf{\Gamma}$ GAUC
Self GPT-2 GPT-2 98.64
Pythia Pythia 96.92
Qwen2.5 Qwen2.5 99.24
Generation Qwen2.5 Qwen2 93.27
Qwen2.5 Qwen1.5 96.10
Qwen2 Qwen1.5 94.21
Family GPT-2 Pythia 92.00
GPT-2 Qwen2.5 91.11
Pythia Qwen2.5 87.39
Figure 7: Relative differences of parameters (blue) and graph matching accuracy (orange) of Pythia-160M model at different checkpoints.
The graph matching extension enables a direct application in LLM fingerprinting (Xu et al., 2024; Russinovich & Salem, 2024), which is crucial for protecting intellectual property. To test this, we perform graph matching on the Pythia-160M model, using checkpoints at 1k, 2k, 4k, 8k, and 16k steps of continued training from a base checkpoint (127k). Figure 7 illustrates the graph matching accuracy measured by GAUC, as well as the relative differences of weight values. Despite significant parameter drift, where the weight difference after 16k steps is 5.57 times greater than after 1k steps, the topological signature remains nearly unchanged, with our method achieving a graph matching GAUC above 0.96 in all cases. This suggests that neural topology can serve as a robust fingerprint, resilient to weight modifications from finetuning, which we propose as a significant avenue for future work.
@@ 6 Related Work
Probing LLMs. Growing concerns over the transparency and steerability of LLMs have driven recent advances in reverse-engineering LLMs by extracting interpretable features from their neural activations through probes (Sharkey et al., 2025). Supervised probing typically maps neuron activations to interpretable semantics through regression or classification (Gurnee et al., 2023; Gurnee & Tegmark, 2024; Jin & Rinard, 2024; Ju et al., 2024; Dong et al., 2023; Kissane et al., 2024; Templeton et al., 2024; Belinkov, 2022). For example, Gurnee et al. (Gurnee & Tegmark, 2024) predicted the time and location of input entities from LLM activations. Unsupervised probing, by contrast, aims to learn a dictionary of disentangled features related to more abstract concepts (Engels et al., 2025; Gao et al., 2025; Rajamanoharan et al., 2024; Lieberum et al., 2024; Mudide et al., 2025; Engels et al., 2025). A famous example is the Golden Gate Bridge feature identified in the Claude 3 Sonnet model (Templeton et al., 2024). While prior work focused on connecting LLM activations to external semantics, our work studies the functional topology of neurons in LLMs, and relates this internal structure directly to language generation performance via graph probing.
Network Neuroscience. The study of functional networks in the human brain has been a central topic in neuroscience for decades (Bassett & Bullmore, 2006; Bassett & Sporns, 2017; Fotiadis et al., 2024; Medaglia et al., 2015) which motivates this research. Brain networks are typically constructed by correlating fMRI or EEG signals across different neural regions, and then analyzed using tools from network science (Barabási, 2013), which has revealed a range of structural and functional properties, such as small-worldness (Bassett & Bullmore, 2006), economical wiring (Bullmore & Sporns, 2012), and functional specialization (Fotiadis et al., 2024). More recently, several studies have drawn parallels between LLM activations and human brain activity (Toneva & Wehbe, 2019; Caucheteux et al., 2023; Kumar et al., 2024; Rathi et al., 2025; Mischler et al., 2024; Tuckute et al., 2024; Bonnasse-Gahot & Pallier, 2024; Sun et al., 2024; Liu et al., 2025). For instance, Tuckute et al. (Tuckute et al., 2024) used GPT-2 activations to identify sentence stimuli that drive or suppress human brain responses. However, while these efforts focus on representational similarities, the functional topology of neurons within LLMs and its relationship to the model’s language generation capabilities remain largely unexplored.
@@ 7 Discussion
Neurons in LLMs are connected both structurally through the model’s architecture and functionally through their dynamic responses to input linguistic stimuli. In this work, we focus on the latter and demonstrate that the language generation performance of LLMs can be reliably predicted from their functional neural topologies using our proposed graph probing approach, implying that LLMs develop intricate and consistent topological structures among their neurons that are fundamental to their emergent ability to generate coherent language. Besides causal intervention on benchmarks validating LLMs actually leverage their internal neural topology, we also offer practical applications of neural topology in model pruning, hallucination detection, and LLM fingerprinting. While we have empirically shown a stable default neural topology and hub neurons in LLMs regardless of input, we have not yet identified more nuanced structures such as motifs, or physical metrics like small-worldness and modularity within these graphs. It remains an open question whether such properties exist in LLMs’ neural topology and play a causal role in shaping their intelligence. Additionally, this paper evaluates LLMs with up to 14B parameters, while leaving graph probing on even larger models for future work.
Our graph probing results raise many interesting directions for future research. While we have linked neural topology to general next-token prediction ability, and discovered functional specialization and subject-specific neurons (Appendix A.5), it remains unclear how these specialized structures emerge during LLMs’ learning process. Additionally, recent advances in enhancing LLMs’ reasoning abilities (Guo et al., 2025) raise a natural question: does reasoning alter, or is it constrained by, neural topology? Finally, graph probing is model-agnostic and can be extended to models other than LLMs, for example to vision-language models (VLMs). We provide preliminary results probing neural topology of VLMs in Appendix A.7-A.8 with causal intervention as well, while further efforts are required to achieve deeper insights into their multi-modal understanding and generation capabilities. In all, we believe graph probing offers a promising lens for understanding AI models and ultimately guiding their improvement in an reliable and safe way.
@@ A.1 Graph Probing Configuration
Hyperparameters. We train graph probes using the Adam optimizer (Kingma & Ba, 2015) with mean squared error (MSE) loss, as defined in Equation (7). The initial learning rate is set to 0.00001, with a batch size of 16. We set the hidden dimension for MLP probes $d$ as 32. We apply a learning rate decay strategy, reducing the rate by a factor of 0.1 if the loss does not improve for 5 consecutive epochs. Each model is trained for up to 100 epochs, with early stopping triggered if no improvement is observed for 20 epochs. Dropout is not used, as preliminary experiments showed no significant impact on regression performance.
@@ A.4 Graph neural network probe
To reduce the number of probe parameters, we adopt a graph neural network (GNN)-based probe that encodes each node by aggregating neighborhood information through convolutional message passing on the graph (Kipf & Welling, 2017; Fey & Lenssen, 2019). We employ the ReLU activation function (Fukushima, 1975) between graph convolution layers and use average and maximum pooling to summarize node-level embeddings into a graph-level representation. Given a connectivity matrix $\mathbf{A}$ induced by feeding a tokenized sequence $X$ to an LLM, where each element $a_{ij}$ denotes the functional connectivity (Pearson correlation coefficient) between neurons $i$ and $j$, our probe produces the graph representation $\mathbf{z}$ as follows:
@@
$\mathbf{z}=\mathtt{AVG\_POOLING}\{\mathbf{\Phi}^{L}_{1,:},\ldots,\mathbf{\Phi}^{L}_{n,:}\}\parallel\mathtt{MAX\_POOLING}\{\mathbf{\Phi}^{L}_{1,:},\ldots,\mathbf{\Phi}^{L}_{n,:}\},$ (16)
@@
Beyond the intrinsic default network, we investigated whether LLMs form specialized networks for different knowledge domains, a plausible consequence of their multi-disciplinary pretraining. Such task-invoked networks have been observed in human brains (Cole et al., 2014). To test this, we leveraged the MMLU-Pro benchmark (Wang et al., 2024) to see if a query’s subject can be predicted from its corresponding neural topology. We framed this as a 10-way classification problem using the 10 subjects with the most samples. For computational efficiency, we applied a simple linear probe to the flattened adjacency matrix, optimized with a cross-entropy loss, similar to Section 5.2. The results of Qwen2.5-0.5B model, shown in Figure 8(a), reveal that neural topology is highly indicative of the subject matter. The topology-based linear probe outperformed a baseline probe on neural activations by 9.39% on average. This advantage held true for nearly every individual subject, with the maximum performance gap exceeding 28.16%. These findings strongly suggest that LLMs develop distinct and linearly separable topological patterns for different knowledge domains, allowing for easy extraction of the context or subject being processed.
@@
A.6 Graph Matching Metrics
Given the predicted similarity matrix $\mathcal{S}\in\mathbb{R}^{N\times N}$ and the target similarity matrix $\mathcal{T}=\mathtt{IDENTITY}(N)$, we calculate the following metrics for graph matching (Li et al., 2019):
$\mathtt{GAUC}=\frac{1}{2N}\sum_{i=1}^{N}{(\mathtt{AREA\_UNDER\_ROC}(\mathcal{S}_{i,:},\mathcal{T}_{i,:})+\mathtt{AREA\_UNDER\_ROC}(\mathcal{S}_{:,i},\mathcal{T}_{:,i}))}.$ (18)

### v3

@@ 4 Topological analysis and causal intervention
To determine if LLMs rely on their neural topology, we conducted a series of interventional experiments. We disabled a selected 1% of neurons in the middle layer of Qwen2.5 models (0.5B, 1.5B, 3B, 7B, 14B) by pinning their activations to zero at all tokens and measured the impact on the MMLU benchmark (Hendrycks et al., 2021). We compared three distinct neuron selection strategies: random selection, selecting neurons with the highest activation, and selecting neurons with the highest topological degree. As shown in the Figure 4(c), the results demonstrate a clear hierarchy of neuronal importance. While disabling random neurons caused a negligible accuracy drop (<1.0%), targeting neurons by either activation or topology incurred a substantial performance collapse of at least 20%. Critically, the topology-based intervention was the most detrimental. Disabling the top 1% of hub neurons resulted in a catastrophic accuracy drop of 45.4%, 57.5%, 61.6%, 66.6%, and 69.8% for the five tested models, consistently outperforming the activation-based strategy across all scales. Particularly, for the largest 14B model, the relative performance drop against activation based intervention is over 95.5%. These experiments provide causal evidence that LLMs actively utilize their underlying topological structure, particularly their hub neurons, for computation.
Our analysis also revealed a notable phenomenon of functional specialization (Cole et al., 2014), where neural topology displays distinct patterns for different subjects. We found a striking example by identifying certain neuron that consistently specializes in STEM subjects, while remaining almost entirely dormant for social science subjects. This suggests a subject-based organization within LLMs, and the details are provided in Appendix A.5.
@@ 5 Applications
Figure 5: (a) Accuracy on MMLU benchmark under different levels of model pruning based on neural topology and activation (WANDA). (b) Accuracy of hallucination detection of different approaches on TruthfulQA dataset. (c) Coupling index of neural topology for hallucination on TruthfulQA datset.
@@
Model pruning. Model pruning is a standard technique for optimizing LLMs in real-world deployments that prioritize low latency and computational efficiency (Sun et al., 2024b; Ma et al., 2023; Gao et al., 2024; Muralidharan et al., 2024; Xia et al., 2024; Xu et al., 2024b). The finding that high-degree neurons are functionally dominant (Section 4) naturally suggests a pruning strategy of disabling low-degree neurons. To test this, we prune these neurons by zeroing out their activations during inference. We apply this technique to a middle layer of Qwen2.5-0.5B and Qwen2.5-1.5B and evaluate the performance on the MMLU benchmark (Hendrycks et al., 2021). We compare our topology based pruning approach with a state-of-the-art LLM pruning method WANDA (Sun et al., 2024b) which is based on neural activation. The results, presented in Figure 5(a), confirm the models’ robustness. Performance degradation is minimal even with substantial pruning; for the 1.5B model, accuracy drops by just 2.86% when 10% of neurons are pruned, and by 11.73% when 25% are removed. Strikingly, the models maintain above-random-guess accuracy until half of the low-degree neurons are pruned. In addition, our topology based pruning approach consistently outperforms WANDA, particularly in the 1.5B model where our method retains over 51% accuracy compared to WANDA’s 30% accuracy when pruning 25% of the neurons. These results highlight the potential for developing more sophisticated pruning methods based on neural topology.
@@
Hallucination detection. Building on our finding that subtle topological patterns in LLMs can predict text generation performance, we now investigate if these patterns can also detect hallucination, a critical challenge that has been extensively explored through various probing techniques (Burns et al., 2023; Du et al., 2024; Hou et al., 2025; Su et al., 2024; Sriramanan et al., 2024; Orgad et al., 2025; Farquhar et al., 2024; Niu et al., 2025). To create distinct genuine and hallucinating states, we construct inputs from the TruthfulQA dataset (Lin et al., 2022) by concatenating each of its 817 question with its corresponding true and false answers, respectively, resulting in a dataset of 5918 samples (true/false classification). We then extract the neural topology from the LLM as it processes these inputs (Equations 1-4). For this task, we adapt probing from regression to binary classification by modifying the MLP probe layer to have two output channels ($\mathbf{W}_{1}\in\mathbb{R}^{2\times(n\times n)}$ and $\mathbf{W}_{3}\in\mathbb{R}^{2\times d}$ in Equations (5-6)) and replacing the MSE loss with a cross-entropy loss:
$\mathcal{L}(\mathbf{X})=$ $\frac{1}{N}\sum_{i=1}^{N}\mathtt{CROSS\_ENTROPY}(\hat{y}_{i},y_{i}),$ (9)
@@
where $\hat{y}$ and $y$ are the prediction and ground-truth for hallucination (0/1). We compare against MLP probes trained on the neural activation at the last-token position. We also include the state-of-the-art HaMI (Niu et al., 2025) and CCS (Burns et al., 2023) hallucination detection approach, as well as CCS’s variant CRC (Burns et al., 2023) for comparison.
@@
We split the dataset into training and test sets using an 8:2 ratio to evaluate our hallucination detection framework. As shown in Figure 5(b), probing neural topology substantially outperforms all baselines probing neural activation for all three tested models (GPT-2, Pythia-160M, and Qwen2.5-0.5B), with accuracy gains of up to 9.73%. This superior performance suggests that distinct topological patterns emerge when an LLM is generating a factual response versus hallucinating. To validate this hypothesis directly, we introduce a neural topology coupling index,
$C_{XY}=\mathtt{AVG}(\{\rho(A_{i},A_{j})|A_{i}\in A_{X},A_{j}\in A_{Y}\}),$ (10)
@@
It is worthwhile to emphasize that our goal with such a simple topological probe is not to establish a new state-of-the-art against more nuanced approaches specifically design for either model pruning or hallucination detection, which are highly specialized domains with extensive literature (Sun et al., 2024b; Ma et al., 2023; Gao et al., 2024; Muralidharan et al., 2024; Xia et al., 2024; Xu et al., 2024b; Kuhn et al., 2023; Niu et al., 2025; Burns et al., 2023; Du et al., 2024; Farquhar et al., 2024; Sriramanan et al., 2024). Rather, these experiments serve as proof-of-concept demonstrations. Though preliminary in their current forms, these case studies showcase the versatile utility of graph probing and lay the groundwork for future work to translate these concepts into robust, system-level solutions. We hope these results encourage the community to view graph probing as a valuable tool for optimizing and interpreting LLMs.
@@ 6 Related Work
Probing LLMs. Growing concerns over the transparency and steerability of LLMs have driven recent advances in reverse-engineering LLMs by extracting interpretable features from their neural activations through probes (Sharkey et al., 2025; Alain and Bengio, 2017; Rogers et al., 2020; Hewitt and Liang, 2019; Voita and Titov, 2020; Pimentel et al., 2020). Supervised probing typically maps neuron activations to interpretable semantics through regression or classification (Gurnee et al., 2023; Gurnee and Tegmark, 2024; Jin and Rinard, 2024; Ju et al., 2024; Dong et al., 2023; Kissane et al., 2024; Templeton et al., 2024; Belinkov, 2022). For example, Gurnee et al. (Gurnee and Tegmark, 2024) predicted the time and location of input entities from LLM activations. Unsupervised probing, by contrast, aims to learn a dictionary of disentangled features related to more abstract concepts (Engels et al., 2025; Gao et al., 2025; Rajamanoharan et al., 2024; Lieberum et al., 2024; Mudide et al., 2025; Engels et al., 2025). A famous example is the Golden Gate Bridge feature identified in the Claude 3 Sonnet model (Templeton et al., 2024). While prior work focused on connecting LLM activations to external semantics, our work studies the functional topology of neurons in LLMs, and relates this internal structure directly to language generation performance via graph probing.
Network Neuroscience. The study of functional networks in the human brain has been a central topic in neuroscience for decades (Bassett and Bullmore, 2006; Bassett and Sporns, 2017; Fotiadis et al., 2024; Medaglia et al., 2015) which motivates this research. Brain networks are typically constructed by correlating fMRI or EEG signals across different neural regions, and then analyzed using tools from network science (Barabási, 2013), which has revealed a range of structural and functional properties, such as small-worldness (Bassett and Bullmore, 2006), economical wiring (Bullmore and Sporns, 2012), and functional specialization (Fotiadis et al., 2024). More recently, several studies have drawn parallels between LLM activations and human brain activity (Toneva and Wehbe, 2019; Caucheteux et al., 2023; Kumar et al., 2024; Rathi et al., 2025; Mischler et al., 2024; Tuckute et al., 2024; Bonnasse-Gahot and Pallier, 2024; Sun et al., 2024a; Liu et al., 2025). For instance, Tuckute et al. (Tuckute et al., 2024) used GPT-2 activations to identify sentence stimuli that drive or suppress human brain responses. However, while these efforts focus on representational similarities, the functional topology of neurons within LLMs and its relationship to the model’s language generation capabilities remain largely unexplored.
@@ 7 Discussion
Neurons in LLMs are connected both structurally through the model’s architecture and functionally through their dynamic responses to input linguistic stimuli. In this work, we focus on the latter and demonstrate that the language understanding and generation performance of LLMs can be reliably predicted from their functional neural topologies using graph probing, implying that LLMs develop intricate and consistent topological structures among their neurons that are fundamental to their emergent linguistic ability. Besides causal intervention on benchmarks validating LLMs actually leverage their internal neural topology, we also offer practical applications of neural topology in model pruning and hallucination detection. While we have empirically shown a stable default neural topology and hub neurons in LLMs regardless of input, we have not yet identified more nuanced structures such as motifs, or physical metrics like small-worldness and modularity within these graphs. It remains an open question whether such properties exist in LLMs’ neural topology and play a causal role in shaping their intelligence. Additionally, this paper evaluates LLMs with up to 14B parameters due to computational cost, while leaving even larger models for future work.
Graph probing raises many interesting directions for future research. While we have linked neural topology to general linguistic ability, and discovered functional specialization and subject-specific neurons (Appendix A.5), it remains unclear how these specialized structures emerge during LLMs’ learning process. Additionally, recent advances in enhancing LLMs’ reasoning abilities (Guo et al., 2025) raise a natural question: does reasoning alter, or is it constrained by, neural topology? Finally, graph probing is model-agnostic and can be extended to LLMs of distinct architectures or even models other than LLMs, for example to vision-language models (VLMs). We provide preliminary results of graph probing for cross-LLM matching and LLM fingerprinting, as well as probing VLMs in Appendix A.6-A.8 with causal intervention as well, while further efforts are required to achieve deeper insights into their multi-modal understanding and generation capabilities. In all, we believe graph probing offers a promising lens for understanding AI models and ultimately guiding their improvement in an reliable and safe way.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning by improving our understanding of the internal mechanisms of large language models. Our work relies exclusively on publicly available, pre-trained models (e.g., GPT-2, Pythia, Qwen2.5) and standard academic benchmarks (e.g., MMLU, TruthfulQA, OpenWebText) to ensure reproducibility and avoid the use of sensitive or private data. We believe our findings have several positive societal implications. The methods for hallucination detection directly contribute to the development of more reliable, truthful, and safe AI systems. Similarly, our work on model pruning promotes computational efficiency, which can reduce the environmental impact of AI and make powerful models more accessible.
@@ A.1 Graph Probing Configuration
Hyperparameters. We train graph probes using the Adam optimizer (Kingma and Ba, 2015) with mean squared error (MSE) loss, as defined in Equation (8). The initial learning rate is set to 0.00001, with a batch size of 16. We set the hidden dimension for MLP probes $d$ as 32. We apply a learning rate decay strategy, reducing the rate by a factor of 0.1 if the loss does not improve for 5 consecutive epochs. Each model is trained for up to 100 epochs, with early stopping triggered if no improvement is observed for 20 epochs. Dropout is not used, as preliminary experiments showed no significant impact on regression performance.
@@ A.4 Graph neural network probe
To reduce the number of probe parameters, we adopt a graph neural network (GNN)-based probe that encodes each node by aggregating neighborhood information through convolutional message passing on the graph (Kipf and Welling, 2017; Fey and Lenssen, 2019). We employ the ReLU activation function (Fukushima, 1975) between graph convolution layers and use average and maximum pooling to summarize node-level embeddings into a graph-level representation. Given a connectivity matrix $\mathbf{A}$ induced by feeding a tokenized sequence $X$ to an LLM, where each element $a_{ij}$ denotes the functional connectivity (Pearson correlation coefficient) between neurons $i$ and $j$, our probe produces the graph representation $\mathbf{z}$ as follows:
@@
$\mathbf{z}=\mathtt{AVG}\{\mathbf{\Phi}^{L}_{1,:},\ldots,\mathbf{\Phi}^{L}_{n,:}\}\parallel\mathtt{MAX}\{\mathbf{\Phi}^{L}_{1,:},\ldots,\mathbf{\Phi}^{L}_{n,:}\},$ (14)
@@
Beyond the intrinsic default network, we investigated whether LLMs form specialized networks for different knowledge domains, a plausible consequence of their multi-disciplinary pretraining. Such task-invoked networks have been observed in human brains (Cole et al., 2014). To test this, we leveraged the MMLU-Pro benchmark (Wang et al., 2024) to see if a query’s subject can be predicted from its corresponding neural topology. We framed this as a 10-way classification problem using the 10 subjects with the most samples. For computational efficiency, we applied a simple linear probe to the flattened adjacency matrix, optimized with a cross-entropy loss, similar to our hallucination detection setup. The results of Qwen2.5-0.5B model, shown in Figure 6(a), reveal that neural topology is highly indicative of the subject matter. The topology-based linear probe outperformed a baseline probe on neural activations by 9.39% on average. This advantage held true for nearly every individual subject, with the maximum performance gap exceeding 28.16%. These findings strongly suggest that LLMs develop distinct and linearly separable topological patterns for different knowledge domains, allowing for easy extraction of the context or subject being processed.
@@
A.6 Model matching and fingerprinting
In this section, we further investigate potential topological similarity across different LLMs to answer the following question: can we detect the genetic relationships of LLMs (e.g. same model family, or finetuning) from their internal neural topology? To this end, we extend graph probing with contrastive learning to perform graph matching, as illustrated in Figure 7. Specifically, suppose we feed a batch of $B$ token sequences into two LLMs, $\Omega$ and $\Gamma$. We compute the corresponding neural connectivity graphs and use two GNN probes to encode them into representations $\mathbf{Z}_{\Omega}=[\mathbf{z}^{\Omega}_{1},\ldots,\mathbf{z}^{\Omega}_{B}]$ and $\mathbf{Z}_{\Gamma}=[\mathbf{z}^{\Gamma}_{1},\ldots,\mathbf{z}^{\Gamma}_{B}]$. Matching is implemented using a contrastive cross-entropy (CE) loss that encourages alignment between graph representations from the same input texts:
$\mathcal{L}=\sum_{i=1}^{B}{\mathtt{CE}(\mathcal{S}_{i,:},\mathcal{T}_{i,:})}+\sum_{j=1}^{B}{\mathtt{CE}(\mathcal{S}_{:,j},\mathcal{T}_{:,j})},$ (17)
where $\mathcal{S}$ is the similarity matrix by taking inner product of graph representations and $\mathtt{IDENTITY}$ indicates identity matrix which is the target $\mathcal{T}$ for graph matching. After training the graph probes contrastively on a shared set of training texts, the out-of-sample graph matching performance serves as an indicator of neural topology similarity between two LLMs. To evaluate this, we adopt the commonly used GAUC metric (Li et al., 2019). Specifically, given the predicted similarity matrix $\mathcal{S}\in\mathbb{R}^{N\times N}$ and the target similarity matrix $\mathcal{T}=\mathtt{IDENTITY}(N)$, GAUC for graph matching is calculated as follow:
$\mathtt{GAUC}=\frac{1}{2N}\sum_{i=1}^{N}{(\mathtt{AUC}(\mathcal{S}_{i,:},\mathcal{T}_{i,:})+\mathtt{AUC}(\mathcal{S}_{:,i},\mathcal{T}_{:,i}))}.$ (18)
Matching LLM $\mathbf{\Omega}$ LLM $\mathbf{\Gamma}$ GAUC
Self GPT-2 GPT-2 98.64
Pythia Pythia 96.92
Qwen2.5 Qwen2.5 99.24
Generation Qwen2.5 Qwen2 93.27
Qwen2.5 Qwen1.5 96.10
Qwen2 Qwen1.5 94.21
Family GPT-2 Pythia 92.00
GPT-2 Qwen2.5 91.11
Pythia Qwen2.5 87.39
Figure 8: Relative differences of parameters (blue) and graph matching accuracy (orange) of Pythia-160M at different checkpoints.
The graph matching extension enables a direct application in LLM fingerprinting (Xu et al., 2024a; Russinovich and Salem, 2024), which is crucial for protecting intellectual property. To test this, we perform graph matching on the Pythia-160M model, using checkpoints at 1k, 2k, 4k, 8k, and 16k steps of continued training from a base checkpoint (127k). Figure 8 illustrates the graph matching accuracy measured by GAUC, as well as the relative differences of weight values. Despite significant parameter drift, where the weight difference after 16k steps is 5.57 times greater than after 1k steps, the topological signature remains nearly unchanged, with our method achieving a graph matching GAUC above 0.96 in all cases. This suggests that neural topology can serve as a robust fingerprint, resilient to weight modifications from finetuning, which we propose as a significant avenue for future work.